A farmer needs to make a triangular fence for his cattle. He decides to make his fence into an isosceles triangle. If the angle that is different in 93°, what will the other two angles be?
Found 2 solutions by macston, edjones:
Answer by macston(5194)

Sum of angle in triangle=180 degrees
S=angles that are the same; D= angle that is different=93 degrees
In an isosceles triangle two angles are the same and one is different.
S+S+D=180 degrees
2S+93 degrees=180 degrees Subtract 93 degrees from each side.
2S=87 degrees Divide each side by 2.
S=43.5 degrees ANSWER: The two similar angles would each be 43.5 degrees.

Answer by edjones(8007)

let a=one of the 2 equal angles
a=(180-93)/2
=87/2
=43.5 degrees
.
Ed